# Effect of symmetry energy on properties of rapidly rotating neutron stars and universal relations

### v1

@@ II Rapidly rotating stars
Fig. 4 shows the dependence of the $f_{\mathrm{Kep}}$ on the kinetic-to-gravitational energy ratio $T/W$ for an illustrative choice of 4 EoS (left panel) and the full set of EoS (right panel). This dependence is useful for locating the dynamical bar-mode instability for $T/W\geq 0.27$ and secular (dissipation driven) bar-mode instability for $T/W\geq 0.14$ for uniformly rotating CSs. The Newtonian results for $T/W$ do not provide useful scaling for our case. For example, for incompressible Maclauring spheriods $T/W={\Omega^{2}R^{3}}/{5GM}$, see Ref. Chandrasekhar1969 , Chapt. 4. Substituting here Eq. (1) shows that, in the Newtonian limit, there is no explicit dependence of $T/W$ on mass or radius. Therefore, we reply below on the numerical results shown in the left panel of Fig. 4 to draw the following conclusions:
  • When all other EoS parameters are held fixed and $T/W$ is constant, a smaller symmetry energy slope $L_{\text{sym }}$ leads to a higher Keplerian frequency $f_{\mathrm{Kep}}$ in the low $T/W\lesssim 0.1$ regime. However, this trend reverses at higher $T/W$, near and beyond the secular instability limit of 0.14.
@@
The trends stated above are also seen in the right panels of Figs. 3 and  4 for the full collection of the EoS, which demonstrates the range of the possible value of Kepler frequency for a broad range of the pair of parameters $30\leq L_{\text{sym }}\leq 110$ MeV and $-600\leq Q_{\text{sat }}\leq 1000$ MeV.
@@ III Universal relations
Figure 5: Log-log plot of the dimensionless moment of inertia $I^{>}$ versus compactness $C$ for maximally fast rotating (Keplerian) configurations. The compactness is defined via the equatorial radius. It is limited from above the maximum mass configuration (the curves are extended slightly in the unstable region where masses are below the maximum). The compactness is limited from below by the configuraltion with mass 0.1 $M_{\odot}$ for $L=30$ MeV and 0.5 $M_{\odot}$ for $L=110$ MeV. The bottom panel shows the relative residual errors $\delta I^{>}$, where $\delta I^{>}=|({I}^{>}-{I}^{>}_{\rm fit})|/{I}_{\rm fit}^{>}$ with respect to the fit by Eq. (3). In each sub-image, the upper panel shows the data points by varying the parameter $L_{\text{sym }}=30,...110$ MeV with $10$ Mev for fixed $Q_{\text{sat }}$ and the best-fit curve based on a third-order polynomial in $C$. The lower panel displays the relative deviation between the numerical data and the fit.
@@
While it is well-established that the macroscopic properties of CSs are highly sensitive to the underlying EoS, several universal relations have been identified between certain global quantities. Here, universality refers to remarkably weak dependence on the EoS, an observation that holds particularly well for cold, $\beta$-equilibrated CSs in the static limit. While the fundamental origin of these universal relations remains not fully understood, they have proven useful in practice. For example, they can be used to constrain otherwise inaccessible stellar parameters, reduce EoS-related uncertainties in data analysis, and break degeneracies between integral observables—such as those between the quadrupole moment and spin in gravitational waveforms from binary inspirals. In this section, we investigate the extent to which these universal relations persist in the case of cold, rapidly rotating CSs, for a set of EoS that exhibit continuous variations in the symmetry energy slope parameter $L_{\text{sym}}$ and the isoscalar skewness coefficient $Q_{\text{sat}}$. Here we focus on relations involving the normalized moment of inertia, the quadrupole moment, and the stellar compactness.
The theoretical minimum mass of a non-rotating neutron star lies in the range 0.2– 0.9 $M_{\odot}$, depending on the EoS. However, neutron stars with masses $M\lesssim 1.17~{}M_{\odot}$ are generally expected to be unstable and unlikely to form via stellar collapse. Instead, such low-mass supernova remnants may result in the formation of a white dwarf.
Rapidly rotating neutron stars, particularly those near the Keplerian (mass-shedding) limit, can be stable at lower masses compared to their non-rotating counterparts. Theoretical estimates place the minimum mass for such configurations in the range 0.1–0.5 $M_{\odot}$, again depending on the EoS. However, the astrophysical formation channels for these extremely low-mass, rapidly rotating neutron stars remain unclear. In the following, we consider sequences of Keplerian (mass-shedding limit) neutron star models with masses $M\geq 0.5\,M_{\odot}$. To properly fit the high-mass end of these sequences, we also include a small portion of the unstable branch (beyond the maximum mass) where the stellar models have masses slightly below the maximum (by a few percent).
We now turn to the two formulas that relate the moment of inertia to the compactness $C=M/R$ where we define the compactness of rotating CS using the equatorial radius. For non-rotating stars the compactness varies in the range $0.1\lesssim C\lesssim 0.3$. For rapidly rotating stars, both the mass and the equatorial radius are increasing, but the equatorial radius increase overwhelms, therefore the compactness of stars close to the maximum effectively decreases, compared to non-rotating neutron stars.
@@
which was proven to be highly EoS-insensitive. Ref. Breu:2016 revisited this problem showing that an alternative universal relation using different normalization of the moment of inertia, $I^{>}=I/M^{3}$ relates to compactness through inverse power expansion:
@@
$\alpha_{0}$ $-6.46939$$\pm 0.04184$ (0.6467%)
@@
$\alpha_{3}$ $-4.52455\times 10^{-4}$$\pm 2.33\times 10^{-6}$ (0.515%)
@@
$\alpha_{5}$ $-1.84774\times 10^{-9}$$\pm 1.293\times 10^{-11}$ (0.6999%)
Table 1: The parameters of the fit for $I^{>}(C)$ function, Eq. (3), after convergence is obtained following 13 iterations. The rms residuals is $\sigma=2.36535$ and the variance of residuals (reduced chisquare) $\chi_{\nu}^{2}=5.59489$.
@@
$\alpha_{0}$ $-2.97357\times 10^{6}$$\pm 1.771\times 10^{5}$ (5.957%)
$\alpha_{1}$ $-2.89435\times 10^{6}$$\pm 1.723\times 10^{5}$ (5.951%)
$\alpha_{2}$ $-1.36802\times 10^{6}$$\pm 8.125\times 10^{4}$ (5.939%)
$\alpha_{3}$ $-40268$$\pm 2.382\times 10^{4}$ (5.915%)
$\alpha_{4}$ $-76473.7$$\pm 4414$ (5.772%)
$\alpha_{5}$ $3.05279\times 10^{6}$$\pm 1.82\times 10^{5}$ (5.962%)
$\alpha_{6}$ $-79219.5$$\pm 4877$ (6.156%)
Table 2: The parameters of the fit for $I^{<}(C)$ function, Eq. (2), after convergence is obtained following 11 iterations with $\sigma=6.55706\times 10^{-3}$ and $\chi_{\nu}^{2}=4.2995\times 10^{-5}$.
Equation (3) was fitted to the full data spanning $Q_{\text{sat }}$ and $L_{\text{sym }}$ using $m=5$. Equation (2) was modified by adding two additional exponential terms that improve the fit in the low $C\leq 0.2$ regime and the polynomial was truncated at $m=4$
Figure 7: Same as in Fig. 5 but for the function $Q^{>}(C)$ fitted using formula (6) for maximally fast rotating (Keplerian) configurations.
Figure 8: Same as in Fig. 5 but for the function $Q^{>}(I^{>})$ fitted using formula (7) for maximally fast rotating (Keplerian) configurations.
The results are shown in Figs. 5 and 6. The upper limit on $C$ is set by the maximum mass configuration. In practice, we include stars beyond the maximum with masses that are below the maximum by a few percent. The lower limit for $C$ corresponds to stars in the mass range $0.1\leq M/M_{\odot}\leq 0.5$; the lower limit here correspondss to $L=30$ MeV while the upper limit to $L=110$ MeV. To enhance the clarity of the exposition, each panel in these figures shows the best fit performed on the full data but only the data for fixed $Q_{\text{sat }}$ is shown in each panel. The results demonstrate that the proposed functional relations remain valid across the collection of EoS under rapid rotation. Tables 1 and 2 list the values of the parameters for Eqs. (3) and (2), respectively.
Ref. Yagi:2013a demonstrated strong universal behavior in the relation between the normalized quadrupole moment and compactness. Defining
$Q\equiv-\frac{Q^{\star}}{M^{3}\chi^{2}}$ (5)
where $Q^{\star}$ is the dimensional stellar (mass) quadrupole moment and $\chi=J/M^{2}$ is the dimensionless spin. They found that $Q$ is nearly a polynomial function of the compactness, $C=M/R$, with a very weak dependence on the EoS, i.e., ${Q}^{<}(C)=\sum_{j=0}^{3}\alpha_{j}C^{j},$ where the coefficients were empirically determined to high precision. Here we will use an alternative that relates the dimensionless quadrupole moment, Eq. (5), to the compactness via Raduta_MNRAS_2020
${Q}^{>}=\sum_{0}^{m}\alpha_{j}[C^{-j}].$ (6)
Fits were performed using formula (6) $m=4$ and the results are shown in Fig. 7.
Final set of parameters Asymptotic Standard Error
$\alpha_{0}$ $-3.90142$$\pm 0.01767$ (0.453%)
@@
$\alpha_{2}$ $-4.21307\times 10^{-3}$$\pm 4.808\times 10^{-5}$ (1.141%)
@@
Table 3: The parameters of the fit for $Q^{>}(C)$ function, Eq. (6), after convergence is obtained following 7 iterations with $\sigma=1.20683$ and $\chi_{\nu}^{2}=1.45644$.
Let us now examine the validity of the $I$-Love-$Q$ universal relations Yagi:2013a ; Yagi:2013b for rapidly rotating CSs, focusing specifically on the ${I}^{>}-{Q}^{>}$ relation. For this purpose, we consider a collection of EoS that feature continuous variation in the isoscalar skewness parameter $Q_{\text{sat }}$ and the symmetry energy slope parameter $L_{\text{sym }}$. If the universality persists in this case, it is reasonable to expect that the remaining relations $I^{>}-\Lambda$ and $Q-\Lambda$ will also hold, given the strong correlation among these quantities observed in previous studies. Here $\Lambda=\lambda^{\star}/M^{5}$ is the dimensionless tidal deformability and $\lambda^{\star}$ is its dimensional counterpart.
The corresponding fit formulas were established already in Refs. Yagi:2013a ; Yagi:2013b , which use the dimensionless moment of inertia $I^{>}$, the dimensionless quadrupole moment $Q$ [Eq. (5)] and the dimensionless tidal deformability $\Lambda$. These can be collectively written as a $m$th-order polynomial on a log-log scale
@@
where $\left(Y_{i},X_{i}\right)$ denotes one of the three quantity pairs listed above and $m=4$ in most cases. The coefficients $\alpha_{j}$ are fitted parameters that can be found from fits to data. Although originally derived in the slow-rotation approximation, these universal relations have been shown to remain valid for rapidly rotating CSs, provided the fit parameters are adjusted to account for rotation frequency Doneva:2013 . Given this evidence, we have performed fits to the case where $Y_{i}\equiv I^{>}$ and $X_{i}\equiv Q^{>}$ using Eq. (7) in the Keplerian limit.
The corresponding data was obtained from the full collection of EoS spanning all values $Q_{\text{sat }}$ and $L_{\text{sym }}$. The fit results are shown in Fig. 8 where each panel compares the best fit to the data with a fixed value of $Q_{\text{sat }}$. It is seen that independent of the $Q_{\text{sat }}$ value the results are uniformly accurate across the values of $L_{\text{sym }}$. In this case, the deviations from the fitted curves are typically at a level of a few percent. This is illustrated in the bottom panel of Fig. 8. We conclude that, for these quantities, universality is preserved for maximally rapidly rotating stars featuring different combinations of symmetry energy (represented by $L_{\text{sym }}$) and high-density pressure in the symmetric matter, represented by $Q_{\text{sat }}$. The fit parameters are listed in Table 4.
@@
$\alpha_{0}$ $4.57104$$\pm 8.883\times 10^{-4}$ (0.02638%)
$\alpha_{1}$ $0.318971$$\pm 4.051\times 10^{-4}$ (0.2051%)
$\alpha_{2}$ $-1.07793\times 10^{-2}$$\pm 1.034\times 10^{-4}$ (1.037%)
$\alpha_{3}$ $1.92037\times 10^{-4}$$\pm 1.724\times 10^{-5}$ (11.53%)
$\alpha_{4}$ $2.70451\times 10^{-5}$$\pm 2.304\times 10^{-6}$ (8.521%)
Table 4: The parameters of the fit for $I^{>}(Q^{>})$ function according to (7) after convergence is obtained following 8 iterations with $\sigma=6.28604$ and $\chi_{\nu}^{2}=39.5143$.

### v2

@@ II Rapidly rotating stars
Figure 4 shows the dependence of $f_{\mathrm{Kep}}$ on the kinetic-to-gravitational energy ratio $T/W$ for an illustrative choice of four EoSs (left panel) and the full set of EoSs (right panel). This dependence is useful for locating the dynamical bar-mode instability for $T/W\geq 0.27$ and secular (dissipation-driven) bar-mode instability for $T/W\geq 0.14$ for uniformly rotating CSs. The Newtonian results for $T/W$ do not provide useful scaling for our case. For example, for incompressible Maclaurin spheriods, $T/W={\Omega^{2}R^{3}}/{5GM}$; see Chap. 4 of Ref. Chandrasekhar1969 . Substituting Eq. (2) here shows that, in the Newtonian limit, there is no explicit dependence of $T/W$ on mass or radius. Therefore, we rely below on the numerical results shown in the left panel of Fig. 4 to draw the following conclusions:
  • When all other EoS parameters are held fixed and $T/W$ is constant, a smaller symmetry energy slope $L_{\text{sym }}$ leads to a higher Keplerian frequency $f_{\mathrm{Kep}}$ in the low-$T/W\lesssim 0.1$ regime. However, this trend reverses at higher $T/W$, near and beyond the secular instability limit of 0.14.
@@
The trends stated above are also seen in the right panels of Figs. 3 and  4 for the full collection of the EoS, which demonstrates the range of the possible value of the Kepler frequency for a broad range of the pair of parameters $30\leq L_{\text{sym }}\leq 110$ MeV and $600\leq Q_{\text{sat }}\leq 1000$ MeV.
@@ III Universal relations
Figure 5: Log-log plot of the dimensionless moment of inertia $I^{>}$ versus compactness $C$ for maximally fast rotating (Keplerian) configurations. The compactness is defined via the equatorial radius. It is limited from above by the maximum mass configuration (the curves are extended slightly in the unstable region where masses are below the maximum). The compactness is limited from below by the configuration with mass 0.1 $M_{\odot}$ for $L=30$ MeV and 0.5 $M_{\odot}$ for $L=110$ MeV. The bottom panel shows the relative residual errors $\delta I^{>}$, where $\delta I^{>}=|({I}^{>}-{I}^{>}_{\rm fit})|/{I}_{\rm fit}^{>}$ with respect to the fit by Eq. (4). In each subpanel, the upper panel shows the data points by varying the parameter $L_{\text{sym }}=30$ (red $\square$), 40 (green $\lozenge$), 50 (blue $\triangle$), 60 (violet ++), 70 (orange $\times$), 80 (magneta $\ast$), 90 (brown $\square$), 100 (rose $\lozenge$) and 110 (grey $\triangle$) in MeV for fixed $Q_{\text{sat }}$ and the best-fit curve based on a third-order polynomial in $C$. The lower panel displays the relative deviation between the numerical data and the fit.
@@
While it is well established that the macroscopic properties of CSs are highly sensitive to the underlying EoS, several universal relations have been identified between certain global quantities. Here, universality refers to remarkably weak dependence on the EoS, an observation that holds particularly well for cold, $\beta$-equilibrated CSs in the static limit. While the fundamental origin of these universal relations remains not fully understood, they have proven useful in practice. For example, they can be used to constrain otherwise inaccessible stellar parameters, reduce EoS-related uncertainties in data analysis, and break degeneracies between integral observables, such as those between the quadrupole moment and spin in gravitational waveforms from binary inspirals. In this section, we investigate the extent to which these universal relations persist in the case of cold, rapidly rotating CSs for a set of EoSs that exhibit continuous variations in the symmetry energy slope parameter $L_{\text{sym}}$ and isoscalar skewness coefficient $Q_{\text{sat}}$. Here, we focus on relations involving the normalized moment of inertia, the quadrupole moment, and the stellar compactness.
The theoretical minimum mass of a nonrotating neutron star lies in the range 0.2–0.9 $M_{\odot}$, depending on the EoS. However, neutron stars with masses $M\lesssim 1.17~M_{\odot}$ are generally expected to be unstable and unlikely to form via stellar collapse. Instead, such low-mass supernova remnants may result in the formation of a white dwarf.
Rapidly rotating neutron stars, particularly those near the Keplerian (mass-shedding) limit, can be stable at lower masses compared with their nonrotating counterparts. Theoretical estimates place the minimum mass for such configurations in the range 0.1–0.5 $M_{\odot}$, again depending on the EoS. However, the astrophysical formation channels for these extremely low-mass, rapidly rotating neutron stars remain unclear. In the following, we consider sequences of Keplerian (mass-shedding limit) neutron star models with masses $M\geq 0.5\,M_{\odot}$. To properly fit the high-mass end of these sequences, we also include a small portion of the unstable branch (beyond the maximum mass) where the stellar models have masses slightly below the maximum (by a few percent).
We now turn to the two formulas that relate the moment of inertia to the compactness $C=M/R$ where we define the compactness of rotating CSs using the equatorial radius. For nonrotating stars, the compactness varies in the range $0.1\lesssim C\lesssim 0.3$. For rapidly rotating stars, both the mass and equatorial radius are increasing, but the equatorial radius increase overwhelms; therefore, the compactness of stars close to the maximum effectively decreases, compared with nonrotating neutron stars.
@@
which was proven to be highly EoS insensitive. Reference Breu:2016 revisited this problem, showing that an alternative universal relation using a different normalization of the moment of inertia, $I^{>}=I/M^{3}$, relates to compactness through inverse power expansion:
@@
$\alpha_{0}$ $2.97357\times 10^{6}$ $\pm 1.771\times 10^{5}$ (5.957%)
$\alpha_{1}$ $2.89435\times 10^{6}$ $\pm 1.723\times 10^{5}$ (5.951%)
$\alpha_{2}$ $1.36802\times 10^{6}$ $\pm 8.125\times 10^{4}$ (5.939%)
$\alpha_{3}$ 40268-40268 $\pm 2.382\times 10^{4}$ (5.915%)
$\alpha_{4}$ 76473.7-76473.7 $\pm 4414$ (5.772%)
$\alpha_{5}$ $3.05279\times 10^{6}$ $\pm 1.82\times 10^{5}$ (5.962%)
$\alpha_{6}$ 79219.5-79219.5 $\pm 4877$ (6.156%)
Table 1: Parameters of the fit for the $I^{<}(C)$ function, Eq. (3), after convergence is obtained, following 11 iterations with $\sigma=6.55706\times 10^{-3}$ and $\chi_{\nu}^{2}=4.2995\times 10^{-5}$.
Final set of parameters Asymptotic Standard Error
$\alpha_{0}$ 6.46939-6.46939 $\pm 0.04184$ (0.6467%)
@@
$\alpha_{3}$ $4.52455\times 10^{-4}$ $\pm 2.33\times 10^{-6}$ (0.515%)
@@
$\alpha_{5}$ $1.84774\times 10^{-9}$ $\pm 1.293\times 10^{-11}$ (0.6999%)
Table 2: Parameters of the fit for the $I^{>}(C)$ function, Eq. (4), after convergence is obtained, following 13 iterations. The rms of residuals is $\sigma=2.36535$ and the variance of residuals (reduced chi-square) $\chi_{\nu}^{2}=5.59489$.
Equation (4) was fitted to the full data spanning $Q_{\text{sat }}$ and $L_{\text{sym }}$ using $m=5$. Equation (3) was modified by adding two additional exponential terms that improve the fit in the low-$C\leq 0.2$ regime, and the polynomial was truncated at $m=4$
Figure 7: Same as in Fig. 5 but for the function $Q(C)$ fitted using Eq. (7) for maximally fast rotating (Keplerian) configurations.
Figure 8: Same as in Fig. 5 but for the function $Q(I^{>})$ fitted using Eq. (8) for maximally fast rotating (Keplerian) configurations.
The results are shown in Figs. 5 and 6. The upper limit on $C$ is set by the maximum mass configuration. In practice, we include stars beyond the maximum with masses that are below the maximum by a few percent. The lower limit for $C$ corresponds to stars in the mass range $0.1\leq M/M_{\odot}\leq 0.5$; the lower limit here corresponds to $L=30$ MeV while the upper limit to $L=110$ MeV. To enhance the clarity of the exposition, each panel in these figures shows the best fit performed on the full data, but only the data for fixed $Q_{\text{sat }}$ is shown in each panel. The results demonstrate that the proposed functional relations remain valid across the collection of EoSs under rapid rotation. Tables 1 and 2 list the values of the parameters for Eqs. (3) and (4), respectively.
Reference Yagi:2013a demonstrated strong universal behavior in the relation between the normalized quadrupole moment and compactness. Defining
$Q\equiv-\frac{Q^{*}}{M^{3}\chi^{2}}$ (6)
where $Q^{*}$ is the dimensional stellar (mass) quadrupole moment and $\chi=J/M^{2}$ is the dimensionless spin, they found that $Q$ is nearly a polynomial function of the compactness, $C=M/R$, with a very weak dependence on the EoS, i.e., $Q(C)\simeq\sum_{j=0}^{3}\alpha_{j}C^{j},$ where the coefficients were empirically determined to high precision. Here we use an alternative that relates the dimensionless quadrupole moment, Eq. (6), to the compactness via Raduta_MNRAS_2020
$Q(C)\simeq\sum_{0}^{m}\alpha_{j}[C^{-j}].$ (7)
Fits are performed using Eq. (7) with $m=4$ and the results are shown in Fig. 7. Table 3 lists the values of parameters for Eq. (7).
@@
$\alpha_{0}$ 3.90142-3.90142 $\pm 0.01767$ (0.453%)
@@
$\alpha_{2}$ $4.21307\times 10^{-3}$ $\pm 4.808\times 10^{-5}$ (1.141%)
@@
Table 3: Parameters of the fit for the $Q(C)$ function, Eq. (7), after convergence is obtained, following 7 iterations with $\sigma=1.20683$ and $\chi_{\nu}^{2}=1.45644$.
Let us now examine the validity of the II-Love-$Q$ universal relations Yagi:2013a ; Yagi:2013b for rapidly rotating CSs, focusing specifically on the ${I}^{>}-Q$ relation. For this purpose, we consider a collection of EoSs that feature continuous variation in the isoscalar skewness parameter $Q_{\text{sat }}$ and the symmetry energy slope parameter $L_{\text{sym }}$. If the universality persists in this case, it is reasonable to expect that the remaining relations $I^{>}-\Lambda$ and $Q-\Lambda$ will also hold, given the strong correlation among these quantities observed in previous studies. Here $\Lambda=\lambda^{*}/M^{5}$ is the dimensionless tidal deformability and $\lambda^{*}$ is its dimensional counterpart.
The corresponding fit formulas were established already in Refs. Yagi:2013a ; Yagi:2013b , which used dimensionless moment of inertia $I^{>}$, dimensionless quadrupole moment $Q$ [Eq. (6)], and the dimensionless tidal deformability $\Lambda$. These can be collectively written as an $m$th-order polynomial on a log-log scale
@@
where $\left(Y_{i},X_{i}\right)$ denotes one of the three quantity pairs listed above and $m=4$ in most cases. The coefficients $\alpha_{j}$ are fitted parameters that can be found from fits to data. Although originally derived in the slow-rotation approximation, these universal relations have been shown to remain valid for rapidly rotating CSs, provided the fit parameters are adjusted to account for rotation frequency Doneva:2013 . Given this evidence, we perform fits to the case where $Y_{i}\equiv I^{>}$ and $X_{i}\equiv Q$ using Eq. (8) in the Keplerian limit.
The corresponding data are obtained from the full collection of EoSs spanning all values $Q_{\text{sat }}$ and $L_{\text{sym }}$. The fit results are shown in Fig. 8, where each panel compares the best fit to the data with a fixed value of $Q_{\text{sat }}$. It is seen that independent of the $Q_{\text{sat }}$ value, the results are uniformly accurate across the values of $L_{\text{sym }}$. In this case, the deviations from the fitted curves are typically at $\lesssim 10\%$. This is illustrated in the bottom panel of Fig. 8. We conclude that, for these quantities, universality is preserved for maximally rapidly rotating stars featuring different combinations of symmetry energy (represented by $L_{\text{sym }}$) and high-density pressure in the symmetric matter (represented by $Q_{\text{sat }}$). The fit parameters are listed in Table 4.
@@
$\alpha_{0}$ =1.17189=1.17189 $\pm 0.02158$ (1.842%)
$\alpha_{1}$ =1.7559=1.7559 $\pm 0.03648$ (2.078%)
$\alpha_{2}$ =0.767783=-0.767783 $\pm 0.02085$ (2.716%)
$\alpha_{3}$ =0.237257=0.237257 $\pm 0.004875$ (2.055%)
$\alpha_{4}$ = -0.021993 $\pm 0.0004255$ (1.935%)
Table 4: Parameters of the fit for the $I^{>}(Q)$ function according to (8) after convergence is obtained, following 13 iterations with $\sigma=5.94561$ and $\chi_{\nu}^{2}=35.3503$.
Figure 9: Comparison of the function $Q(I^{>})$ fitted using Eq. (8) for Keplerian CSs (left panel) and CSs in the limit of slow rotation (right panel) The conventions are the same as in Fig. 5.
Given that the II-Love-$Q$ relations exhibit smaller errors in the slow-rotation limit than in the Keplerian limit, we perform the computation in the slow-rotation regime by fixing the frequency to 1 Hz and using a reduced number of members (30) in each sequence defined by $Q_{\text{sat }}$ and $L_{\text{sym }}$. The Keplerian and slow-rotation results are compared in Fig. 9 for the case $Q_{\text{sat }}=0$. As expected, in the slow-rotation case, the relative error is suppressed to just a few percent, with the most significant differences appearing when $Q\leq 10$.